OpenStudy (anonymous):

If c(x) = 4x – 2 and d(x) = x2 + 5x, what is c(d(x))

OpenStudy (anonymous):

hey you again :D

Nnesha (nnesha):

ello same question :=)

OpenStudy (anonymous):

so just do the same thing?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

but x is already x so would it be 2x^2 + 5x^2 in the D function?

Nnesha (nnesha):

that would be the same so just replace x in c(x) function by d(x)

Nnesha (nnesha):

\[\huge\rm c(\color{ReD}{d(x)}) = c(\color{Red}{x^2+5x}) =4\color{reD}{x}-2\]

OpenStudy (anonymous):

so am i just simplifying x^2+5x=4x-2?

Nnesha (nnesha):

no you need to plug x^2+5x into the x c(x^2+5x) is same as c(d(x))

Nnesha (nnesha):

4x -2 ^^^ that x replace by x^2 +5x

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

4x^2 + 3

Nnesha (nnesha):

mhm.. how did you get that ?

OpenStudy (anonymous):

4 and x^2 combine then 5-2

Nnesha (nnesha):

remember when multiply same bases we should add their exponents \[\huge\rm x^m \times x^n = x^{m+n}\] and on't forget the parenthesis

Nnesha (nnesha):

it should be like this \[\huge\rm 4(x^2+5x)-2\] distribute parentheses by 4 and then combine like terms

OpenStudy (anonymous):

ohh i forgot parentheses

OpenStudy (anonymous):

4x2+20x−2??

Nnesha (nnesha):

:=) right